OpenStudy (anonymous):

I need help with substitution. 2m + n = 1 m - n = 8

hartnn (hartnn):

ok, find/isolate m from 2nd equation. m-n=8 add n on both sides, what you get ?

OpenStudy (anonymous):

m = 8 + n?

hartnn (hartnn):

thats correct! now substitute that in 1st equation, what u get ?

OpenStudy (anonymous):

2(8 + n ) + n = 1?

hartnn (hartnn):

yes, now distribute 2 , 2(8+n) =.. ?

OpenStudy (anonymous):

16 + 2n + n = 1 :3

hartnn (hartnn):

going good :) now 2n+n =...? also subtract 16 on both sides.

OpenStudy (anonymous):

16 + 3n = 1. Then subtract 16?

hartnn (hartnn):

yes!

OpenStudy (anonymous):

n = -5?

hartnn (hartnn):

good! its correct :)

OpenStudy (anonymous):

m = 3?

hartnn (hartnn):

correct!

OpenStudy (anonymous):

The solution is going to be (3,-5) right?

hartnn (hartnn):

yes, thats right.

OpenStudy (anonymous):

Mkaii. :3 i have another question but i'll make a new post. c:

hartnn (hartnn):

okay :)